A sphere, a cube and a disc made of same material and of equal massesheated to same temperature of 200°C. These bodies are then kept atsame lower temperature in the surrounding, which of these will cool(i) fastest, (ii) slowest, explain.

Grade:11

1 Answers

Satdhruti Paul
6 years ago
Let their mass be M. Obviously, the one with largest surface area will release heat fastest!
Since they are of same material, density is same. D = M/V so V = M/D
in Sphere 4/3(pi)R^3 = M/D
in Cube a^3 = M/D
in disc (pi)r^2*(thickness) = M/D.
From here you can proceed by finding R, a & r. Find surface area respectively and check.